OpenStudy (anonymous):

How many solutions does this linear system have? y = 2x – 5 –8x – 4y = –20

OpenStudy (anonymous):

A. one solution: (–2.5, 0) B. one solution: (2.5, 0) C. no solution D. infinite number of solutions

OpenStudy (mrnood):

the first equation tells you that you can write y in terms of x (i.e wherever there is y it is the SAME as 2x-5) SO In the second equation you can substitute (2x-5) WHEREVER there is y that will allow oyu to find a solution for x....

OpenStudy (lynfran):

y=2x-5 y=(-20+8x)/-4 \[2x-5=-\frac{ 8x-20 }{ 4 }\]\[8x-20=-8x+20\]\[8x+8x=20+20\]

OpenStudy (lynfran):

what did u get when u solve for x??